### v1

# Replay Consolidation with Label Propagation for Continual Object Detection

Abstract
Object Detection is a highly relevant computer vision problem with many applications such as robotics and autonomous driving. Continual Learning (CL) considers a setting where a model incrementally learns new information while retaining previously acquired knowledge. This is particularly challenging since Deep Learning models tend to catastrophically forget old knowledge while training on new data. In particular, Continual Learning for Object Detection (CLOD) poses additional difficulties compared to CL for Classification. In CLOD, images from previous tasks may contain unknown classes that could reappear labeled in future tasks. These missing annotations cause task interference issues for replay-based approaches. As a result, most works in the literature have focused on distillation-based approaches. However, these approaches are effective only when there is a strong overlap of classes across tasks. To address the issues of current methodologies, we propose a novel technique to solve CLOD called Replay Consolidation with Label Propagation for Object Detection (RCLPOD). Based on the replay method, our solution avoids task interference issues by enhancing the buffer memory samples. Our method is evaluated against existing techniques in CLOD literature, demonstrating its superior performance on established benchmarks like VOC and COCO.
Keywords Continual Learning  $\cdot$ Object Detection
Figure 1: Continual Learning for Object Detection pipeline. The model learns to detect new classes at each incremental training stage - corresponding to a new task. The additional issue is the missing annotations; the labeled example in the current task might have been learned as background in previous tasks (e.g., the class dog is labeled in Task 1 but not in Task 2 or 3.).
1 Introduction
Object Detection is a critical computer vision problem involving both the localization and classification of objects within an image, with applications from robotics to autonomous driving [1]. However, a key challenge of Deep Learning systems is that they cannot learn incrementally over time, as they are prone to Catastrophic Forgetting [2, 3] when training on new data. Continual Learning (CL) addresses this problem, introducing techniques to learn new information over time while retaining previously acquired knowledge.
Many CL studies focus on image classification, where replay-based approaches are the most common way to achieve strong results [4, 5, 6, 7]. These approaches involve storing a subset of samples from previous tasks in a memory buffer and replaying them when training on new data. However, Continual Learning for Object Detection (CLOD) is more complex because images seen in previous tasks may contain unlabeled classes that the model must learn as new classes in future tasks, causing missing annotations and training conflicts [8]. For example, the image in Fig. 1 has the classes person, dog, and car. However, this image has only one labeled class for each task. Then, this image could be saved multiple times in the replay memory, each time with different annotations based on the task, causing interference. Because of this, most of the CLOD literature is based on distillation-based approaches [8]. However, these approaches encounter several challenges, particularly the tendency to forget older classes when their instances are underrepresented in new training images.
To address the limitations of replay approaches for CLOD, we propose a novel technique called Replay Consolidation with Label Propagation for Object Detection (RCLPOD). Our approach consists of multiple components, including a technique called Label Propagation [9], a selection mechanism called Optimizing Class Distribution in Memory (OCDM) [10], a Feature Distillation technique [11], and a Masking Loss [12]. The key idea of our approach is to significantly reduce the task interference issue of the replay memory via the Label Propagation technique, enabling more efficient use of the samples saved in the replay memory. The superior performance of our approach is validated on the well-known CLOD benchmarks based on VOC and COCO [8].
While in the literature some CL studies were conducted either on YOLOv3 [12] and v5 [13], we test our approach on YOLOv8 [14], a recent and anchor-free version of the real-time object detector YOLO [15]. Therefore, this work can serve as a valuable benchmark for future research on contemporary YOLO architectures within the CLOD framework.
Overall, we make the following contributions:
  • We propose a novel approach called RCLPOD, which addresses the main issues of replay techniques in the CLOD setting.
  • We conduct extensive experiments on RCLPOD, demonstrating its superior performance on the established VOC and COCO CLOD benchmarks.
  • We benchmark our approach and other CL techniques using more modern versions of YOLO compared to other works in the literature.
The paper is structured as follows. Sec. 2 provides an overview of the CL literature focusing on CLOD. In Sec. 3, we describe our novel approach, RCLPOD. Sec. 4 describes the experimental setting, including the considered metrics, evaluated scenarios, and experimental details. In Sec. 5, we show and discuss the results. Finally, Sec. 6 concludes this work by discussing limitations and future research directions.
2 Related Work
Continual Learning aims to enable DL models to continuously accumulate knowledge over time without the need to re-train from scratch and with limited access to past data [16]. In particular, this framework has emerged to tackle the Catastrophic Forgetting problem [2, 3]; after a DL model is trained on one task and subsequently trained on another, it tends to forget the previously learned task. The methods in the CL literature can be grouped into three big families: rehearsal, regularization, and architecture-based approaches.
Most of the literature on Continual Learning for Object Detection focuses on the Class Incremental Learning (CIL) scenario, where each new task introduces new classes [8]. For example, Fig. 1 depicts a CLOD setting composed of three tasks where a new class must be learned each time. At the end of the training, the model should be able to provide predictions on all these classes. Moreover, most of the works were conducted on two-stage object detectors, whose time complexity scales linearly with the number of predicted objects in the scene. On the contrary, one-stage object detectors are known to be able to perform real-time object detection. In particular, recent papers focus on anchor-free architectures, with the most known probably being the modern versions of YOLO [15, 14, 17].
The first CL approach proposed for a CLOD scenario work is  [18]. The authors propose to use the Learning without Forgetting (LwF) regularization-based approach in the context of a two-stage object detector, Fast-R-CNN [19]. LwF [20], following the idea of knowledge distillation [21], forces the outputs of the training model, called student, to be similar to the corresponding outputs of a frozen version of the model trained on old tasks, called teacher. Following this work, many subsequent approaches in the literature are distillation-based approaches [8].
However, all of them consider either two-stage object detectors [19, 22] or old one-stage object detectors like YOLOv3 [23], which are not anchor-free like FCOS [24] or more recent versions of YOLO like v8 and v10 [17, 14].
More recent works applied to anchor-free fully convolutional detectors [25] revisit regularization-based algorithms for the FCOS object detector. Indeed, this type of model has a separate branch for regression outputs that introduces noise when training with techniques like LwF. Building on this work, a solution for CLOD on the Edge called Latent Distillation was proposed [26]. However, these solutions are tailored FCOS-based architectures, which makes its adoption in modern YOLO architectures challenging.
Another popular regularization-based approach is Pseudo Label [27]. Before training on the new task, the new task samples are passed through the old model, and its predicted labels are concatenated to the ground truth. However, like all distillation-based approaches, the main issue with pseudo-label is that performance remains high until images containing classes of the previous tasks are present in the new images. Otherwise, performance decreases quickly.
Also, Replay is used in the CLOD literature[8]. This method stores a portion of the old data in a memory buffer. While training on a new task, the new data is combined with the data from the memory buffer to maintain knowledge of old tasks. According to the CL literature for image classification, Replay is one the most effective solutions for catastrophic forgetting [4, 28].
However, this is not the case in object detection, where task interference arises caused by the missing annotations. When training on a new task, images saved in the memory buffer during previous tasks may also present unlabeled instances of new classes. This means that for those buffer samples, the ground truth interferes with the new class knowledge that the model is trying to learn.
To address these issues, the authors of [12] building on the YOLOv3 architecture proposed to mask the loss associated with new classes during the training of old samples and vice-versa for the new data. However, the masking loss only partially solves the problem caused by the missing annotations. Indeed, this approach fails to fully leverage the potential information contained within the replay memory. Indeed, when saving the current task training samples in the replay memory, the images are labeled only on the classes of the current task but might contain instances of new classes (example in Fig. 3), leading to an under-utilization of the replay memory’s potential.
(a) RCLPOD approach during the training of a new task.
(b) Consolidation of the information inside the replay memory using the Label Propagation technique.
Figure 2:
3 Replay Consolidation with Label Propagation for Object Detection
To tackle the complexities of Object Detection within a Continual Learning framework, we combine multiple components to maximize their effectiveness and achieve optimal results.
The first component of our method is a Replay memory, but instead of randomly selecting the images for the memory buffer, we adopt the OCDM selection mechanism [10], which aims to have a memory buffer with a more balanced class distribution (see Fig. 4(b)). This selection mechanism is advantageous because it is tailored for settings like CLOD, where each image is associated with multiple labels. The second component is the Label Propagation mechanism, which fully leverages the memory buffer potential by integrating old and new knowledge into its ground truths, thus addressing the interference problems inherent in Replay. To further reduce the task interference problem, we introduce a Masking Loss specifically designed for the anchor-free YOLOv8 model, as explained in Sec. 3.2. Finally, to further decrease the forgetting of old knowledge, we also apply the Feature Distillation technique [11] as presented in Sec. 3.3.
3.1 Label Propagation
In multi-label classification and object detection, replay-based methods commonly face interference issues. This interference arises between the current task samples and the replayed samples when there is an overlap between task classes, as each sample ground truth only has information pertinent to the labels of the associated task. Moreover, the full potential of Replay buffers is not realized in these scenarios, as each sample is stored in the memory buffer with ground truth labels limited to the specific classes of its original task (see Fig. 3(a)).
Label Propagation, illustrated in Fig. 2, is composed of the forward and backward steps.
In the forward step, information from old labels is added to the new data during training and subsequently into the replay buffer through a pseudo-labeling technique. In detail, let $D_{t}=\{X_{t},Y_{t}\}$ be the data of the current task $t$ with labels corresponding to a set of classes $C_{t}$. During the training of task $t$, for a sample $(x,y)\in D_{t}$, the forward step modifies the ground truth $y$ adding pseudo-labels of the classes $C_{1},\ldots,C_{t-1}$ by leveraging the knowledge of the previously trained model $f_{\theta_{t-1}}$. Indeed, the model $f_{\theta_{t-1}}$ provides predictions associated with classes $C_{1},\ldots,C_{t-1}$.
(a) Replay Memory
(b) RCLPOD Memory
Figure 3: Comparison of the memory for Replay and for RCLP. The vertical axis represents the stored samples of each task, while the horizontal axis represents the objects associated with each class $c_{i}$. (a) Each task of the Replay memory has information only on the classes seen during its iteration. (b) Using the Label Propagation mechanism of RCLP, the saved samples are more informative, containing knowledge of new and old classes.
(a) Replay Memory
(b) OCDM Memory
Figure 4: Example showing the difference between Replay and OCDM by comparing the class distribution of the memory buffer. The vertical axis represents the label frequency, while the horizontal axis represents the different labels. (a) On the left, the distribution of Replay is unbalanced. (b) At right, the distribution of OCDM is more balanced because of the selection mechanism.
After training on task $t$, a subset of the samples from task $t$ is stored in the memory buffer $M$. This method ensures that, after training on task $t$, the memory buffer retains samples with information relevant to all tasks up to and including task $t$. On the other hand, the backward step focuses on consolidating the new knowledge gained from training the model on the latest task into the old samples stored in the replay buffer. Specifically, using the new model trained on task $t$, pseudo-labels corresponding to the current task are added to the ground truths of old samples of the memory buffer.
3.2 Masking Loss
Figure 5: Image from replay memory: class "tennis racket" is a new class, while class "person" is an old one. Since the only ground truth available is the one for the "person" class, any model prediction for the tennis racket would be penalized in the classification loss computation.
Label Propagation considerably reduces task interference in replay-based techniques. However, the issue of handling new classes for replay memory samples when training on new tasks remains. On the other hand, since YOLO uses Task Alignment Learning (TAL) [29] for label-prediction assignment, the extent to which interference affects the performance depends on the saved images. In fact, for a given image and its corresponding ground truths, the YOLO assigner assigns at least one model prediction to each ground truth based on the intersection area between the ground truths and the predictions. Consequently, if an object of a new class is not close to an object of an old class, the prediction for the new class object is not assigned to any ground truth, meaning that this prediction doesn’t contribute to the loss.
Therefore, depending on the specific dataset, if instances of new-class objects don’t overlap frequently with those from previous tasks, interference does not arise; otherwise, as in Fig. 5 where, without any class masking, the tennis racket prediction would be penalized in the loss computation, samples from the replay memory can prevent the model from learning new classes effectively.
To tackle this problem independently of the dataset, we propose a masking loss approach similar to the method used in [12]. In this method, the contribution of new classes for computing the classification loss is ignored for the samples saved in the memory buffer; recalling that, for a prediction assigned to a ground truth, the classification loss is given by $\mathcal{L}_{\text{cls}}=\sum_{i}\,\mathcal{L}_{\text{cls}}^{i}$, where $\mathcal{L}_{\text{cls}}^{i}$ is the usual Binary Cross Entropy for class $i$, by applying the masking approach, the classification loss for any sample in the Replay memory is given by:
$\mathcal{L}_{\text{cls-mask}}=\sum_{i\not\in C_{t}}\,\mathcal{L}_{\text{cls}}^{i}$ (1)
where $C_{t}$ is the set of the new classes at task $t$. In the case of samples for the current task, the classification loss is the original one instead.
3.3 Feature Distillation
Finally, to further enhance the method’s performance, we employ the Feature Distillation technique [11], similar to RCLP and SID approaches[25]. The core idea behind feature distillation is that the features generated by the new model for both new and old data should remain close to those produced by the previous model for the same data. This approach helps the model maintain a more cohesive memory, mitigating the effects of catastrophic forgetting and ensuring more stable performance across tasks.
In our method, we perform feature distillation on the model’s intermediate representations. In details, let our model $f(x)=g_{\phi}(h_{\omega}(x))$, where given an input $x$, $h$ gives the intermediate features, and $g$ produces the model output based on such features. Formally, given an input sample $x$, the distillation loss on intermediate features is given by:
$\mathcal{L}_{\text{feat\_dist}}=||h_{\omega_{t}}(x)-h_{\omega_{t-1}}(x)||_{2}^{2}$ (2)
where $\omega_{t}$ denotes the current model parameters while $\omega_{t-1}$ the ones belonging to the old model trained on the previous task.
However, the YOLOv8 architecture includes a neck in addition to the usual backbone $h_{\omega}$. Therefore, we propose to compute the distillation loss on the intermediate features of the backbone and the neck. Recalling that the YOLOv8 backbone outputs three feature maps at multiple levels $C_{3},C_{4},C_{5}$ and the neck does the same $P_{3},P_{4},P_{5}$ [30, 31], the overall loss can be computed as:
$\mathcal{L}=\mathcal{L}_{\text{YOLO}}+\lambda\mathcal{L}_{\text{feat\_dist}}$ (3)
where $\mathcal{L}_{\text{YOLO}}$ is the YOLOv8 loss, while $\mathcal{L}_{\text{feat\_dist}}$ is the feature distillation loss. Specifically, for YOLOv8, we consider the distillation of 6 feature maps $C_{3},C_{4},C_{5},P_{3},P_{4},P_{5}$, 3 of the necks and 3 of the backbone, rewriting the generic equation 2 for YOLO as follows:
$\mathcal{L}_{\text{feat\_dist}}=\sum_{i=3}^{6}||C_{i}^{\omega_{t}}-C_{i}^{\omega_{t-1}}||_{2}^{2}+||P_{i}^{\omega_{t}}-P_{i}^{\omega_{t-1}}||_{2}^{2}$ (4)
3.4 OCDM
The OCDM method was originally proposed to work in the multi-label image classification setting [32]. Its main contribution is proposing a new selection mechanism that differs from the classic random selection used when constructing the replay memory. Referring to Fig. 4, OCDM addresses the unbalanced label distribution found in replay memory samples, which makes them unrepresentative for many non-frequent labels. This is done by ensuring a more uniform label frequency distribution via a greedy approach that removes samples one by one based on how much they contribute to the distribution. We integrate OCDM’s selection mechanism into RCLPOD to determine which samples to include in the replay memory after it has been enhanced by the Label Propagation technique.
4 Experimental Setting
4.1 Scenarios and Metrics
[…]
As in [18] to evaluate the performance of YOLOv8, we report the mean average precision (mAP) at a 0.5 IoU threshold (mAP50) for VOC and the mAP weighted across different IoU from 0.5 to 0.95 (mAP50-95) for COCO.
4.2 YOLO training details
To evaluate how the CL performances are affected by the model size, in our study we consider two different versions of YOLOv8: YOLOv8n with 3.2M parameters and YOLOv8m with 25.9M parameters. We initialize the backbone parameters with the ones pre-trained on Image-Net, available on [14].
In Table 3 we report the hyper-parameters used for training. In particular, by following [14, 17], we use SGD with Nestorov momentum and weight decay. For both the learning rate and the momentum we have 3 warmup epochs, while just for the learning rate we employ a linear decay scheduling from $10^{-2}$ to $10^{-4}$. In all the experiments we set the number of epochs per task to 100, which is a suitable value for reaching convergence.
4.3 CL strategies details
We compare the following methods against RCLPOD: Joint training, Fine-tuning, Replay, OCDM, LwF, and Pseudo-Label. In particular, while Replay, LwF, and Pseudo-label were already evaluated in the CLOD setting, this is the first time, to the best of our knowledge, that OCDM approach is implemented and tested in the CLOD setting. Join Training is considered as an upper bound and corresponds to the model performances when trained on the entire dataset with all the classes. Fine-tuning is used as a lower bound; each task is trained sequentially, and no CL technique is employed to avoid catastrophic forgetting.
Regarding the Replay methods, the memory has a small capacity fixed to around 5% of the entire dataset, namely 800 images for PASCAL VOC and 6,000 for COCO. Once the model is trained on task $i$, $n$ samples are randomly selected from the current training set and added to the Replay memory. To prevent exceeding the memory’s total capacity when adding $n$ new samples, $n$ samples are randomly removed from the replay memory. For OCDM, we use the same memory size as all the other replay-based approaches.
For LwF, by following [18], we set the distillation loss weight $\lambda$ to 1. For Pseudo-Labeling, to ensure consistency with inference, we set the classification threshold to $0.5$ and the IoU threshold, for Non-maximum Suppression, to $0.7$. For RCLPOD we use the same hyper-parameters of Pseudo-Labels and Replay. As for LwF, we set the gain for the feature distillation loss to $\lambda=1$.
Scenarios
VOC (mAP 50) COCO (mAP 50-95)
CL Method 10p10 19p1 15p5 15p1 40p40 40p10
Joint Training 78.5 78.5 78.5 78.5 37.3 37.3
Fine-Tuning 36.4 16.4 26.5 2.9 13.6 2.9
Replay 54.5 60.9 56.9 40.0 14.1 8.0
LwF 36.7 3.7 13.5 3.1 8.8 1.9
OCDM 56.5 59.9 54.7 50.2 16.7 11.0
Pseudo-label 71.4 46.6 59.7 16.3 24.3 16.1
RCLPOD (ours) 72.5 68.4 67.7 58.3 24.1 20.0
Table 1: Results for YOLOv8n. Each column represents one of the studied scenarios based on VOC and COCO datasets. Each row represents a different tested CL technique. In bold the best method and underlined the second best method. As in CLOD literature, mAP50 is used for VOC-based scenarios and mAP 50-95 for COCO-based scenarios.
5 Results
In this section, we discuss the outcomes of each method on the VOC and COCO CLOD benchmarks. Table 1 reports the results after all tasks have been completed for YOLOv8n.
In Sec. 5.1, we discuss the results obtained in the 2-tasks scenarios, namely 15p5, 10p10, 19p1, and 40p40. In Sec. 5.2, we consider the more challenging scenarios 15p1 and 40p10.
5.1 Two tasks scenarios
The results for all the two task scenarios are reported in Tab. 1. Specifically, scenarios 10p10, 19p1, and 15p5 for the VOC dataset and scenario 40p40 for the COCO dataset.
As expected, using only distillation, as in LwF, does not guarantee good performances for YOLOv8, as it performs even worse than fine-tuning. Indeed, as discussed in Sec. 2, the noisy regression outputs of the teacher prevent the model from learning the new classes and avoiding forgetting. Therefore, a simple L2 loss between the teacher and the student outputs is highly incompatible with the YOLO training.
The results for the Replay method show strong performance on the VOC dataset, particularly in the 10p10 and 19p1 scenarios, where it achieves mAP scores of 54.5 and 60.9, respectively. This indicates that Replay can effectively retain knowledge from earlier tasks when applied to simpler datasets like VOC. When comparing the Replay method with OCDM, we can observe similar results except for the scenario COCO 40p40, which has a slight improvement.
Notably, it performs better than Replay in the 15p1 scenario with a mAP of 50.2, highlighting its effectiveness in handling challenging incremental tasks.
Regarding Pseudo-label, we can observe its effectiveness for most of the scenarios except for the 19p1, where the gap with respect to RCLPOD is significant. We suppose this is due to the low number of old-class objects in the images for the new tasks; the lower the number of classes for the second task, the less overlap between tasks we have (see the plots of Fig. 9 in the Appendix). On the contrary, this is not the case for the 40p40 scenario, where Pseudo-label, despite its simplicity, reaches the best mAP. Even if the overall mAP is higher than the RCLPOD one, a higher forgetting is observed, while plasticity brings that higher final result. For more details, we refer to Sec. .2 Appendix, where we compare RCPLOD and Pseudo-label, the two best methods in the 40p40 scenario, in the case of the bigger model YOLOv8m.
(a) Scenario VOC 15p1 with the mAP50
(b) Scenario COCO 40p10 with mAP 50-95.
Figure 6: Results obtained for VOC 15p1 at left and COCO 40p10 at right. Ours method RCLPOD performs best in both scenarios.
5.2 Scenarios VOC 15p1 and COCO 40p10
Here, we discussed the two longer and more challenging streams: VOC 15p1 and COCO 40p10. As mentioned earlier, COCO 40p10 is more complex since the VOC dataset is much smaller. However, the analysis of VOC 15p1 is still meaningful because the overlap in images among tasks is much smaller in VOC than in COCO.
The first observation is that LwF doesn’t perform well with the YOLOv8 architecture for COCO and VOC, achieving comparable performance at Fine-Tuning (lower bound). This confirms the results obtained for the simpler and shorter scenarios.
In contrast to these approaches, Pseudo-Label works very well in a complex scenario like COCO 40p10. This is partially due to the very high intersection between tasks in terms of images, which allows the model to maintain knowledge of previous classes as it implicitly revisits them in the new task. However, in cases where the intersection between tasks is very low the Pseudo-Label performance is expected to be much worse. This is demonstrated by the VOC 15p1 scenario, where performance is low as only one new class is shown, and the intersection with previous tasks is much lower compared to COCO 40p10, where ten new classes are added each time.
Therefore, in practice, we have this interesting duality aspect where Replay for VOC 15p1 performs better than Pseudo-Label and vice versa for the COCO 40p10 scenario.
Unlike Pseudo-label, interference between tasks in Replay is a problem since it implies saving the same image in memory several times with different labels, causing several issues, including an under-exploitation of memory.
OCDM, despite having the same problem of missing annotations, is rewarded by carefully selecting the samples to save, which improves the final performance from 8 for Replay to 11 for OCDM in the COCO 40p10.
Then RCLPOD, which exploits OCDM’s selection mechanism and the Label Propagation technique, receives a significant boost, achieving 20 in the COCO 40p10. This proves how much help the correct management of the information of the samples stored in the replay memory can be effective.
5.3 Ablation study
Replay OCDM LP FD Mask mAP
Model 1 14.1
Model 2 16.7
Model 3 19.7
Model 4 20.1
Model 5 24.3
Model 6 22.7
Table 2: Ablation study performed to study the effect of each component in RCLPOD, tested on COCO 40p10. LP is the Label Propagation mechanism, OCDM indicates the selection mechanism, FD represents the Feature Distillation component, and Mask represents the Masking Loss component.
Here, we discuss the contribution of each component to the final performance of our novel approach, RCLPOD. Recall that the components are OCDM, Label Propagation (LP), Feature Distillation (FD), and Masking Loss (Mask). Tab. 2 reports the ablation studied, where the effect of each component is studied.
First, we compare Replay (Model 1) with OCDM (Model 2). Even in a short CL scenario like COCO 40p40, OCDM performs better than Replay, showing the positive effect of balancing the Replay memory. By adding the Label Propagation (LP) mechanism (Model 3), a significant improvement of 3 mAP is obtained. In fact, by doing so, we both reduce the interference problem, and we fully utilize the potential of the replay memory. Feature Distillation (Model 4) brings an additional improvement of 0.4 mAP. Once we add the Masking Loss component (Model 5), namely, we use all the components of RCLPOD, we get a further increase of 4.1 mAP. In fact, as shown in Fig.10, the important overlap of many classes between tasks brings interference. Therefore, when a Replay memory is employed, masking is needed even in the case of modern versions of YOLO. Finally, we test OCDM once all the other components are used. By removing OCDM (Model 6), a 1.6 mAP drop is observed, showing the importance of balancing the replay memory in reducing forgetting, particularly for longer streams such as VOC 15p1 and COCO 40p10.
6 Conclusions and Future Work
To solve the Continual Object Detection, we propose a novel approach called RCLPOD, composed of multiple components such as the Label Propagation technique to exploit, at best, the replay memory, the Feature Distillation to maintain better the old representations, and using a selection mechanism fitted than the classic random mechanism of Replay. By effectively composing these components for the YOLOv8 architecture, we can obtain excellent results for the CLOD setting. We used well-known datasets, such as VOC and COCO, to prove the superiority of RCLPOD to the other methods tested in CLOD.
While our method has allowed us to reduce the gap in the literature with upper-bound Joint Training, it is clear that further work is required to achieve performance comparable to Joint training. An issue that our current approach does not address is that while the introduction of pseudo-labels in replay memory has proven to be beneficial, the noise of such labels must be taken into account as well. Indeed, it is realistic to expect the model to generate false positives and negatives, which will be kept in the replay memory, partially causing noise to the model.
Moreover, we believe that future research should give more attention to longer streams like COCO40p10 and beyond to better represent the real performances of the CL techniques in the CLOD setting.
Appendix
.1 Further training hyper-parameters details
Table 3 presents the hyper-parameters used for training YOLO. Regarding data augmentation, we follow [14, 17].
Table 3: Hyper-parameters for YOLOv8
Hyper-parameter
epochs $100$
optimizer SGD
batch size $8$
momentum $0.937$
weight decay $5\times 10^{-4}$
warm-up epochs $3$
warm-up momentum $0.8$
warm-up bias learning rate $0.1$
initial learning rate $10^{-2}$
final learning rate $10^{-4}$
learning rate scheduling linear decay
box loss gain $7.5$
classification loss gain $0.5$
DFL loss gain $1.5$
.2 Results with YOLOv8m
CL Method YOLOv8n YOLOv8m
Task 1 old new all Task 1 old new all
Pseudo-label 34 22.7 26.0 24.3 47.1 39.2 37.5 38.4
RCLPOD 25.7 22.5 24.1 40.9 36.4 38.6
Table 4: Results for YOLOv8n and YOLOv8m in the COCO 40p40 scenario. "old" is the mAP for the classes of Task 1 after Task 2 training, "new" is the mAP for the new classes, while "all" is the mAP over all the 80 classes.
As shown in Tab. 1, Pseudo-label performs slightly better than RCLPOD. However, as shown in 4, Pseudo-label performs worse than RCLPOD in terms of forgetting, while it outperforms the latter by plasticity.
As proved in [35], robustness to forgetting improves with scale of model size. Therefore, we also investigate the effect of the model size in the COCO 40p40 scenario. As presented in 4, RCLPOD performs better than Pseudo-Label, and the gap in terms of plasticity is lower than in the case of YOLOv8n. This fact shows that once the model size increases, the intrinsic plasticity of the model allows one to acquire new knowledge independently by the CL strategy applied to prevent forgetting.
.3 Label Distribution for each VOC and COCO scenario
Figure 7 to Figure 12 show the class frequencies for each scenario and for each task. In particular, for a given task $t$, in orange we have the classes for task $t$, while in blue, the classes for either old tasks or new ones. These plots are useful to analyze the possible interference between tasks. Moreover, these plots show why Pseudo-label is performing poorly in the scenarios VOC 19p1 and VOC 15p1. For instance, in the 19p1 scenario, many classes for task 1 are missing in the dataset for task 2, namely Pseudo-label is not effective for those classes. This is not the case for replay-based approaches, like RCLPOD. On the other hand, due to the high number of classes, Pseudo-label is more effective in the 40p40 scenario, where instead replay methods struggle due to interference and masking is needed as shown in Sec. 5.
.4 Comparison with literature
In Table 5, we report the results obtained in other CLOD works in the 40p40 COCO scenario. Notice that all of them are specifically designed for different architectures with different loss functions, namely Table 5 doesn’t aim to compare the CL methods but to state the difference of our work based on a more recent Object Detector as YOLOv8. Moreover, we report both the results for YOLOv8n and YOLOv8m, where the latter is comparable in terms of the number of parameters to ResNet50, used as the backbone in the other works.
CL Method mAP 50-95 Model size
ILOD [18] 21.30 $\geq 25$M
IncDet [36] 29.70 $\geq 25$M
Faster ILOD [37] 20.64 $\geq 25$M
SID [25] 25.20 $\geq 25$M
Meta-ILOD [38] 23.80 $\geq 25$M
RCLPOD (YOLOv8n) 24.09 $3.2$M
RCLPOD (YOLOv8m) 38.64 $25.9$M
Table 5: MS COCO Incremental 40p40 literature results.
(a) Task 1
(b) Task 2
Figure 7: Classes distribution over tasks for VOC 10p10. In blue the classes without labels.
(a) Task 1
(b) Task 2
Figure 8: Classes distribution over tasks for VOC 15p5. In blue the classes without labels.
(a) Task 1
(b) Task 2
Figure 9: Classes distribution over tasks for VOC 19p1. In blue the classes without labels.
(a) Task 1
(b) Task 2
Figure 10: Classes distribution over tasks for COCO 40p40. In blue the classes without labels.
Refer to caption Refer to caption Refer to caption
Task 1 Task 2 Task 3
Refer to caption Refer to caption
Task 4 Task 5
Figure 11: Classes distribution over tasks for COCO 40p10. In blue the classes without labels.
(a) Task 1
(b) Task 2
(c) Task 3
(d) Task 4
(e) Task 5
(f) Task 6
Figure 12: Classes distribution over tasks for VOC 15p1. In blue the classes without labels.